OpenStudy (precal):

Site needs to add and equation editor on the ASK part. Many people are posting incorrect problems because they do not use () or ^. This would be solved by adding an equation editor in the ASK part.

OpenStudy (osanseviero):

\[\left( 1\div3 \right)\int\limits_{2}3^{a}{?}\left[\begin{matrix}1 & 2 \\ 3 & 4\end{matrix}\right]\] There is already a good equation editor... there is \[\left( ? \right) and 3^{3} \]

OpenStudy (osanseviero):

There is no real need for putting it in the question, just put the equation as an anwer. (yes, it would make life faster, but not too much)

OpenStudy (precal):

But sometimes, we guess what the person meant to post. Such a waste of time and when people post things incorrectly it is because they are not aware of proper sytax. example x+3/2 but they mean (x+3)/2